The mean test score for a class was 70 points, and the standard deviation was 15 points. If I add 10 points to everyone's test score..
(a) What is the mean test score for the class after I add the 10 points to each score?
(b) What is the standard deviation of the test scores for the class after I add the 10 points to each score?

Answer by Theo(13342)

i believe the mean will be 80 and the standard deviation will remain at 15.

a simple experiment indicates that this will be the case.

assume 3 numbers, for example:

50
100
150

the mean will be 50 + 100 +150 = 300 / 3 = 100

the standard deviation will be (50-100)^2 + (100-100)^2 + (150-100)^3 = 2500 + 0 + 2500 = 5000/3 = 1666 and 2/3

add 10 points to everybody's score.

the scores becomes:

60
110
160

the mean is 60 + 110 + 160 = 330 / 3 = 110

the standard deviation is (60 - 110)^2 + (110-110)^2 + (160-110)^2 = 2500 + 0 + 2500 = 5000 / 3 = 1666 and 2/3.

the mean was raised by 10 and the standard deviation stayed the same.

here's an excel printout that shows the same thing.

the numbers were randomly generated by excel.

then the mean and the standard deviation were calculated for both the original data set and the data set where each element was raised by 10.

the mean was 10 greater and the standard deviation remained the same.

$$$